If the ionization enthalpy and electron gain enthalpy of an element are 275 and 86 kcal respectively then the electronegativity of the element in the Mulliken scale is......

Grade:11

2 Answers

IHate Ppts
6 years ago
Electronegativity = Electron affinity+Ionisation potential/2
Electronegativity=275+86/2
Electronegativity=180.5 (in respective units)
Ritik
4 years ago
Xm=I.P+E.A/125 (In kcal/mol this formula is used bcz energies are given in kcal/mol) 
So ans = 275+86/125 =2.888